OpenStudy (anonymous):

One positive number is one-fifth of another number. The difference between the two numbers is 164, find the numbers. The two numbers in increasing order are _______ and _______ .

OpenStudy (anonymous):

the numbers are x and x/5. x - x/5 = 164 4x/5 = 164 x = 205

OpenStudy (anonymous):

(the other number is 205/5 = 41)